HeartTricker:

A line whose slope is 4 and passes through the point (0, -9) lies on the line whose equation is: A. 4x + y = -9 B. 4x - y = 9 C. y - 4x = 9

AZ:

y = mx + b m is the slope b is the y-intercept so replace 'm' with 4 and then to solve for 'b', plug in the point (0, -9) That's in the form of (x, y) so replace x with 0 and y with -9 and then solve for b

HeartTricker:

can I have the equation so its clear for me

AZ:

Let's do it, step by step The slope-intercept form for an equation is this y = mx + b they told you the slope is 4 replace 'm' with 4 what do you get so far when you replace 'm' with 4

HeartTricker:

\[y= 4x+b \]

AZ:

Good now they told you this line passes through the point (0, -9) points are written in the form of (x, y) so replace x with 0 and y with -9 what do you get then?

HeartTricker:

\[-9=40+b\]

dunlapace:

that wat i about said

AZ:

no, 4x means 4 * x so 4 * 0 -9 = (4 * 0) + b

AZ:

what is 4 * 0 = ? and then can you solve for b?

HeartTricker:

ohh ok

HeartTricker:

0 right?

AZ:

yes, 4 * 0 = 0 so we have -9 = 0 + b so what is b?

HeartTricker:

maybe y? i dont know

AZ:

uhh no? anything plus 0 is going to be that number so b is -9

AZ:

so we have y = mx + b where m is 4 and we just found that 'b' is -9 we have y = 4x - 9 can you re-arrange this equation and see which of the answer choices it would match up to?

AZ:

Hint: subtract y on both sides and then add 9 on both sides